Using the data in the table below, calculate the CPI and the inflation rate in each year using 2005 as the base year.

Instructions: Round your answers to one decimal place.
  

Year Price of basket
($)
CPI Inflation rate
2005 20,000
2006 21,500
2007 22,800
2008 26,150
2009 28,825
2010 32,700
Answer #1

Ans:

CPI = (Cost of basket in current year / Cost of basket in base year) × 100

Inflation rate = % change in CPI.

Calculation of CPI and the inflation rate

Year Price of basket CPI Inflation rate
2005 $20,000 ($20,000 / $20,000) * 100 = 100 -
2006 $21,500 ($21,500 / $20,000) * 100 = 107.5 (107.5 - 100)/100 = 7.5%
2007 $22,800 ($22,800 / $20,000) * 100 = 114 (114 - 107.5)/107.5 = 6.05%
2008 $26,150 ($26,150 / $20,000) * 100 = 130.7 (130.7 - 114)/114 = 14.6%
2009 $28,825 ($28,825 / $20,000) * 100 = 144.1 (144.1 - 130.7)/130.7 = 10.2%
2010 $32,700 ($32,700 / $20,000) * 100 = 163.5 (163.5 - 144.1)/144.1 = 13.5%